# Optimal Mechanisms for Demand Response: An Indifference Set Approach

### v1

@@ Abstract
The time at which renewable (e.g., solar or wind) energy resources produce electricity cannot generally be controlled. In many settings, consumers have some flexibility in their energy consumption needs, and there is growing interest in demand-response programs that leverage this flexibility to shift energy consumption to better match renewable production—thus enabling more efficient utilization of these resources. We study optimal demand response in a model where consumers operate home energy management systems (HEMS) that can compute the “indifference set” of energy-consumption profiles that meet pre-specified consumer objectives, receive demand-response signals from the grid, and control consumer devices within the indifference set. For example, if a consumer asks for the indoor temperature to remain between certain upper and lower bounds, a HEMS could time use of air conditioning or heating to align with high renewable production when possible. Here, we show that while price-based mechanisms do not in general achieve optimal demand response, i.e., dynamic pricing cannot induce HEMS to choose optimal demand consumption profiles within the available indifference sets, pricing is asymptotically optimal in a mean-field limit with a growing number of consumers. Furthermore, we show that large-sample optimal dynamic prices can be efficiently derived via an algorithm that only requires querying HEMS about their planned consumption schedules given different prices. We demonstrate our approach in a grid simulation powered by OpenDSS, and show that it achieves meaningful demand response without creating grid instability.
@@ 1 Introduction
Consumer electricity demand is largely unmanaged.We are grateful to Andrey Bernstein, Mahmoud Saleh and Xinyang Zhou for valuable feedback and numerous helpful conversations; and also thank Xinyang Zhou for helping us get set up with the OpenDSS grid simulator. This research was funded by a grant from Trellis Climate (Prime Coalition) to advance deep demand flexibility solutions and a grant from the Office of Naval Research. In most markets today, consumers obtain electricty via fixed-rate contracts that do not adapt to real-time grid conditions or reflect the true costs of electricity production. This unmanaged demand is in tension with renewable energy resources, such as solar and wind: Unlike conventional energy resources, renewables yield intermittent production and cannot be dispatched on demand, thus leading to frequent mismatches between electricity production and consumption (Gowrisankaran et al., 2016; Cordera et al., 2023). This supply-demand mismatch then often forces the grid to depend on expensive and less efficient power plants during peak demand periods, leading to increased operational expenses and environmental impacts (Bird et al., 2013; Impram et al., 2020; Michaelides, 2019).
@@
As renewable energy sources like solar and wind become more prevalent—for example, in the U.S., electricity demand is expected to grow by approximately 1% annually from 2022 to 2050, with renewable sources projected to supply 44% of electricity by 2050, up from 21% in 2021 (U.S. Energy Information Administration, 2023)—these inefficiencies are projected to intensify further, mainly due to the mismatch between renewable energy availability and peak demand periods. This mismatch is exacerbated by the growing adoption of rooftop solar panels, which allow households to both generate and consume electricity, thereby causing local demand profiles to fluctuate unpredictably. These fluctuations require more sophisticated balancing mechanisms to maintain grid stability Caballero-Pena et al. (2022), highlighting the need for more adaptive and responsive demand management solutions.
To address this challenge, demand-response programs have gained attention by leveraging user flexibility (Yoon et al., 2014; Godina et al., 2018; Siano, 2014). For example, demand response may involve pre-cooling: On a hot day, it may be advantageous to apply air conditioning early in the afternoon while solar production is plentiful, thus enabling reduced electricity consumption during evening peak hours. However, real-world implementations of demand response programs have encountered significant limitations.
[…]
In Texas, a more advanced system has been introduced where consumers are incentivized to lower their energy usage relative to a predetermined baseline. This program, known as the Emergency Response Service, compensates participants for curtailing electricity use during grid emergencies. However, there have been concerns about the program’s fairness and effectiveness due to potential manipulation by strategic consumers (Office of Science and Technology Policy, 2022); for example, some cryptocurrency miners have been accused of inflating their baseline consumption to maximize payouts when reducing usage.
These examples illustrate the challenges in designing demand-response programs that are both effective in managing grid costs and robust against adversarial behaviors, highlighting the need for more sophisticated demand-response management strategies.
1.1 Demand Response with Home Energy Management Systems
Home Energy Management Systems (HEMS) offer a robust solution for automating household energy management by dynamically adjusting consumption plans in response to grid conditions and demand-response signals, demonstrating significant potential for improving demand-response strategies. These systems allow users to set preferences, such as desired indoor temperature ranges, and then operate devices to meet these preferences while optimizing energy use (Beaudin and Zareipour, 2015; Zhou et al., 2016). By interacting directly with the grid, HEMS can pre-cool or pre-heat homes when renewable energy is abundant, reducing demand during peak periods without compromising comfort. This approach enhances grid flexibility and efficiency seamlessly, without placing impractical decision-making burdens on users (Faruqui et al., 2010).
The first fundamental question that arises when implementing demand-response systems with HEMS is: What kind of demand flexibility can the HEMS exploit? In full generality, one could imagine each customer writing down a complete welfare function and then having the HEMS optimize it; and this is the standard starting point in most of the existing literature (Adelman and Uçkun, 2019; Zhou et al., 2017). Consumers have a documented willingness of, e.g., refraining from washing clothes in response to higher prices during peak hours (Allcott, 2011a). Furthermore, this type of HEMS could in principle navigate complex tradeoffs: For example, some consumers may be willing to sell the energy stored in their EV battery (and forgo transportation) when prices are high, while other may be willing to set adaptive indoor temperature targets depending on prices. In practice, however, there’s likely a limit to how well consumers can directly reason about welfare implications of energy use choices in the face of fluctuating energy prices because of to the high required cognitive costs (Friedman and Oren, 1995; Hobman et al., 2016).
In this paper, we instead focus on a narrower form of demand response, which does not seek to shape consumer behavior and instead simply leverages under-specification in the consumers’ existing preferences. Perhaps a consumer wants their EV charged by morning—but doesn’t care exactly when the charging happens; or perhaps a consumer sets upper and lower bounds on a thermostat, but doesn’t care how the indoor temperature evolves between these bounds. We model this underspecification as leading to indifference sets containing a number of equally desirable energy consumption profiles; and consider the demand response problem in which we seek to induce the HEMS to pick consumption profiles from within the indifference set in a way that is well aligned with overall grid objectives.
Given this formulation of the demand response problem, the second fundamental question is: What kind of signals should the grid send to the HEMS to achieve optimal demand response? At one extreme, a decentralized system could be implemented where the grid uses dynamic pricing, and the HEMS simply seek to minimize costs while meeting user preferences. At the other extreme, in a fully centralized system, the HEMS could give the grid transparent control over all devices, allowing the grid to solve a global optimization problem. We refer to these two mechanisms as the (dynamic) pricing mechanism and the direct control mechanism. We study both mechanisms, and seek to understand how close the pricing mechanism can get to direct control in terms of demand response metrics. In particular, we show the following:
@@ 1.1 Demand Response with Home Energy Management Systems
    Under our indifference-set model for demand response, the pricing mechanism described above is generally not optimal relative to the direct control mechanism. A grid operator with direct control of all consumer devices can achieve better demand response (while respecting user indifference sets) that an operator who can only control demand via dynamic prices. [^1] This finding may seem surprising at first glance, as there are numerous welfare theorems in economics demonstrating that price-based mechanisms can achieve optimal welfare in various systems (Mas-Colell et al., 1995). Our indifference-set model, however, is inconsistent with a key assumption of these welfare theorems, often referred to as local non-satiation. Local non-satiation implies that consumer demand moves smoothly with prices—but this is not the case here, as in our setting we only consider hard constraints on the user side.
@@
    Mean-field optimal prices can be recovered via a first-order algorithm that only requires querying HEMS about planned consumption profiles at different price levels, as these consumption profiles give information about a dual cost function. This allows for the straightforward application of first-order optimization methods to find the optimal price using only market-level data, without the grid needing to directly query for individual user preferences.
Finally, we demonstrate practical promise of our pricing approach on a grid-simulation motivated by pre-cooling with data from Phoenix, Arizona; we use OpenDSS to solve the relevant power-flow equations. As seen in Figure 1, renewable production is heavily concentrated during daytime hours. With flat rates, this results in the notorious “duck curve” with low (even negative) net demand during the day that then quickly ramps up in the evening. [^2] One should imagine the ramp-up period as forming the neck of the duck and the evening peak demand its head (Denholm et al., 2015). In contrast, our proposed algorithm shifts demand earlier into the day when solar resources are plentiful, thus enabling it to reduce load in the evening hours. We give further details on this experiment in Section 5.
@@ 1.2 Related work
As the demand for electricity continues to rise and the efficiency and environmental benefits of renewable energy sources grow, there has been significant interest in demand-response programs. Demand response programs are generally classified into two main categories: Dynamic pricing and incentive-based programs. The former use time-varying prices to better align demand with supply (Adelman and Uçkun, 2019; Peura and Bunn, 2015; Uckun, 2012; Crew et al., 1995), while the later involves compensating users through specific contracts to adhere to designated consumption patterns during certain time windows.
@@
Pricing-based mechanisms for demand response are typically framed as Stackelberg games with a two-level leader-follower structure (Adelman and Uçkun, 2019; Zhou et al., 2017, 2019): First the grid sets prices, and then consumers respond to these prices. Our indifference-set approach is a special case of such a Stackelberg games, where lower-level optimization involves consumers choosing their energy consumption to minimize costs while remaining within their indifference sets (i.e., the consumer preferences are instantiated as hard constraints on allowed consumption patterns). We note that Stackleberg games are also often used for the more abstract task of policy learning with strategic agents (Chen et al., 2020; Munro, 2024; Perdomo et al., 2020; Sahoo and Wager, 2022); and our paper can be seen as a part of this broader literature that exploits specific structure of consumer energy price response.
@@
Perhaps closest to us in the literature on price-based mechanisms, Adelman and Uçkun (2019) study dynamic pricing in a setting that includes current energy demand, appliance status, operation times, global variables such as weather conditions, among many others. They assume that each consumer’s welfare depends on their energy usage in a differentiable way, and that consumers choose their energy consumption to optimize welfare in response to set prices (e.g., using a HEMS). Given this setup, the authors define the social planner’s welfare function as the sum of all consumers’ welfare minus the grid’s generation cost, and seek prices that optimize the social planner’s welfare under the posited consumer behavior model (the consumers choose how much to consume in response to prices, and the grid must then meet this demand). Their main result is that, as the system size grows large, the optimal price mechanism in their model is one that perfectly externalizes the grid’s incremental production costs to the consumers, i.e., in equilibrium, energy is priced to match the marginal cost of production. The authors propose finding prices that maximize the social planner’s welfare via a tempered fixed-point iteration algorithm.
@@
Like Adelman and Uçkun (2019), we consider a model where—at least in a qualitative sense—prices are set in the view of mutual benefit of consumers and producers. However, unlike Adelman and Uçkun (2019), we do not rely on properties of the consumer welfare function, and do not posit a social planner who is able to trade-off consumer and producer welfare on an additive scale. Instead, while in our model producers still have a quantitative (differentiable, convex) cost function, consumer welfare is only captured via hard constraints (i.e., we must respect their indifference sets); the main role of prices is then to induce consumers to substitute peak-hour demand with non-peak demand. Several mathematical tools used by Adelman and Uçkun (2019)—such as the implicit function theorem to capture consumer response to prices—are no longer applicable in our setting, and thus our technical development is not closely related to theirs.
@@
In another related line of work, Zhou et al. (2017) consider demand-response programs that employ price incentives, where each consumer aims to minimize their own cost function. This cost function is composed of two parts: one that reflects power consumption and remains independent of price incentives (e.g., minimizing deviation from a nominal consumption level), and another that accounts for the expenses incurred due to price incentives. Building on this framework, Zhou et al. (2019) extends the analysis to incorporate discrete decision variables and dynamically coupled devices over time, such as HVAC systems. Their approach shares similarities with ours, particularly in modeling consumer objectives as minimizing cost functions and solving a social-welfare optimization problem (the direct method in our setup) to determine optimal pricing strategies. However, there are notable differences. In their model, the grid’s utility function is solely governed by voltage constraints, without considering the cost of supply as a function of total consumption. In this setup, and assumptions including strong convexity of consumers’ cost functions (the first component), the social welfare bi-level optimization problem can be reformulated into a convex problem, making it tractable to find optimal prices. These results, however, do not apply to our setting as the consumer cost functions are not strongly convex under our indifference-set approach—and in fact pricing mechanisms generally fail to be efficient in our model.
@@
Meanwhile, on the side of incentive-based programs, Agrawal and Yücel (2022), Daniels and Lobel (2014), and Webb et al. (2019) have explored direct load control contracts, where utility companies manage customer loads, typically to reduce consumption as a form of demand response. Fattahi et al. (2023) formulates a stochastic dynamic program to design optimal contracts and develop approximation methods to achieve near-optimal solutions for energy consumption. They avoid explicitly referring to these contracts as demand response due to their complexity and the extended nature of the control periods. In contrast, in the pricing algorithm we evaluate in our experiments, the grid has no explicit control over users’ consumption; instead, HEMS selects optimal consumption profiles that meet users’ needs while minimizing costs. Finally, Allcott (2011b) studies the effect of “home energy reports” that compare a consumer’s energy use to that of their neighbors, and finds that such programs can meaningfully shift behavior.

### v2

@@ Abstract
The time at which renewable (e.g., solar or wind) energy resources produce electricity cannot generally be controlled. In many settings, however, consumers have some flexibility in their energy consumption needs, and there is growing interest in demand-response programs that leverage this flexibility to shift energy consumption to better match renewable production—thus enabling more efficient utilization of these resources. We study optimal demand response in a setting where consumers use home energy management systems (HEMS) to autonomously adjust their electricity consumption. Our core assumption is that HEMS operationalize flexibility by querying the consumer for their preferences and computing the “indifference set” of all energy consumption profiles that can be used to satisfy these preferences. Then, given an indifference set, HEMS can respond to grid signals while guaranteeing user-defined comfort and functionality; e.g., if a consumer sets a temperature range, a HEMS can precool and preheat to align with peak renewable production, thus improving efficiency without sacrificing comfort. We show that while price-based mechanisms are not generally optimal for demand response, they become asymptotically optimal in large markets under a mean-field limit. Furthermore, we show that optimal dynamic prices can be efficiently computed in large markets by only querying HEMS about their planned consumption under different price signals. Using an OpenDSS-powered grid simulation for Phoenix, Arizona, we demonstrate that our approach enables meaningful demand response without creating grid instability.
@@ 1 Introduction
Consumer electricity demand is largely unmanaged: We are grateful to Andrey Bernstein, Mahmoud Saleh, Jing Shang and Xinyang Zhou for valuable feedback and numerous helpful conversations; and also thank Xinyang Zhou for helping us get set up with the OpenDSS grid simulator. This research was funded by a grant from Trellis Climate (Prime Coalition) to advance deep demand flexibility solutions and a grant from the Office of Naval Research. In most markets today, retail customers obtain electricity via fixed-rate contracts that do not adapt to real-time grid conditions or reflect the true costs of electricity production. [^1] Many regions also run wholesale energy markets where electricity is traded at floating prices between utilities, power producers, municipalities, etc. (Wolak, 2021). End consumers, however, generally do not trade directly on these markets; and exposing consumers to wholesale prices is generally considered to lead to unacceptable risks, as illustrated by the Griddy Energy case below. This unmanaged demand is in tension with renewable energy resources, such as solar and wind: Unlike conventional energy resources, renewables yield intermittent production and cannot be dispatched on demand, thus leading to frequent mismatches between electricity production and consumption (Gowrisankaran et al., 2016; Cordera et al., 2023). This supply-demand mismatch then often forces the grid to depend on expensive and less efficient power plants during peak demand periods, leading to increased operational expenses and environmental impacts (Bird et al., 2013; Impram et al., 2020; Michaelides, 2019).
@@
As renewable energy sources like solar and wind become more prevalent—for example, in the U.S., electricity demand is expected to grow by approximately 1% annually from 2022 to 2050, with renewable sources projected to supply 44% of electricity by 2050, up from 21% in 2021 (U.S. Energy Information Administration, 2023)—these inefficiencies are projected to intensify further, mainly due to the mismatch between renewable energy availability and peak demand periods. This mismatch is further exacerbated by the growing adoption of rooftop solar panels, which allow households to both generate and consume electricity, thereby causing local demand profiles to fluctuate unpredictably. These fluctuations require more sophisticated balancing mechanisms to maintain grid stability (Caballero-Pena et al., 2022), highlighting the need for more adaptive and responsive demand management solutions.
These challenges have led to increasing interest in demand-response programs that seek to leverage consumer flexibility to achieve demand patterns better aligned with renewable production (Yoon et al., 2014; Godina et al., 2018; Siano, 2014). For example, demand response programs may encourage consumers to engage in pre-cooling: On a hot day, it may be advantageous to apply air conditioning early in the afternoon while solar production is plentiful, thus enabling reduced electricity consumption during evening peak hours. Similarly, these programs may incentivize households returning home to shift electric vehicle charging to nighttime hours when demand is lower.
However, real-world implementations of demand response programs have encountered significant limitations in practice. In one of the most high profile applications of demand response, California’s state grid operator has deployed “Flex Alert” text message alerts to urge residents to reduce electricity use during peak demand periods, especially during extreme heat waves. [^2] https://www.latimes.com/california/story/2022-09-07/a-text-asked-millions-of-californians-to-save-energy-they-listened-averting-blackouts While this method has at times effectively prompted immediate public action, it relies heavily on voluntary participation and may not consistently achieve predictable load reductions. In particular, consumers showed evidence of messaging fatigue following repeated use of Flex Alerts in 2022; and not all consumers were able to correctly implement some grid-supporting measures (such as pre-cooling) even when they attempted to comply with a Flex Alert (Opinion Dynamics, 2024). As such, it may be implausible to hope to achieve large-scale, systematic changes to consumer behavior via alert-based methods that rely on consumer goodwill alone.
At the other end of the spectrum, some jurisdictions have explored eliminating fixed-rate energy contracts and instead directly exposing consumers to wholesale energy pricing. For example, in 2017, a company called Griddy started offering consumers in Texas the option to directly purchase energy at wholesale prices. [^3] https://www.caller.com/story/news/local/texas/state-bureau/2021/03/01/griddy-electric-bills-spiked-during-freeze-explanation/4549742001/ However, this system broke down when Texas faced electricity shortages during severe winter storms in 2021: [^4] https://www.nytimes.com/2021/02/20/us/texas-storm-electric-bills.html Wholesale energy prices spiked $\sim$300 times higher than usual and Griddy was forced into bankruptcy as many consumers were unable to pay their bills (Busby et al., 2021). In the aftermath of Griddy’s bankruptcy, consensus formed that exposing consumers to unpredictable price fluctuations of this magnitude was not acceptable, and Texas passed a law that prohibits directly exposing residential customers to wholesale energy prices (Hernandez, 2021).
The Electric Reliability Council of Texas (ERCOT), which oversees the Texan power grid, has also designed a more targeted demand-response program, in which some loads are pre-registered as “Emergency Response Service” (ERS) loads that can be curtailed—in exchange for payment—during energy shortages (Baldick, 2021). Conceptually, these ERS loads can then be understood as “virtual power plants” in the sense that, when the grid to faces an energy shortage, they can choose to either increase generation capacity or pay the ERS to curtail (Mnatsakanyan and Kennedy, 2014). Unfortunately, however, such pay-to-curtail schemes appear in practice to be gameable, especially by market participants that are able to flexibly schedule (and cancel) loads at peak hours. For example, The Economist recently reported that some cryptocurrency companies were able to collect demand-response payments of the same order of magnitude as their overall revenue: [^5] https://www.economist.com/united-states/2024/08/27/why-texas-republicans-are-souring-on-crypto
“On the hottest and coldest days, when demand for electricity peaks and the price rockets, the bitcoin miners sell power back to providers at a profit or stop mining for a fee, paid by ERCOT. Doing so has become more lucrative than mining itself. In August of 2023 Riot [^6] Riot Platforms is a publicly traded cryptocurrency company. collected \$32m from curtailing mining and just \$8.6m from selling bitcoin.”
Incentive-compatibility issues can also arise in more mundane settings. For example, Pacific Gas & Energy in California has experimented with a program called SmartAC which compensates consumers for not using their air conditioners during system-wide or local system strain (Wang et al., 2023); however, it is hard to stop such programs from compensating consumers at times when they never intend to use their air conditioners anyways (e.g., while they are traveling).
These challenges highlight the need for practical mechanisms for demand response that make it financially rational for consumers to respond to demand-response signals, that do not subject consumers to excessive risks or attention burdens, and that are simple and transparently non-gameable. The goal of this paper is to outline such a system, and to establish its basic operating characteristics and optimality properties in settings with many customers. The core of our approach involves modeling consumer flexibility using an “indifference set” approach, whereby each consumer has a pre-determined set of goals [^7] The fact that we take the consumer indifference sets as given reflects the fact that we are not trying to change consumer preferences (e.g., to accept less comfortable indoor temperatures in order to save energy), but rather to leverage flexibility inherent to existing preferences to achieve meaningful demand response. they want to have met (e.g., keep their indoor temperature comfortable, have their electric vehicle charged by morning, etc.), but is indifferent to how these goals are accomplished. We further assume that each consumer has access to a device, called a Home Energy Management System (HEMS), [^8] Devices of this type have been widely discussed in the literature (Beaudin and Zareipour, 2015; Zhou et al., 2016); and in fact many consumers already own smart thermostats whose complexity is comparable to what we require of a HEMS (Stopps and Touchie, 2021), like Google Nest. that can receive price signals from the grid and rationally control devices such as to minimize energy costs subject to meeting the consumer’s goals. [^9] Essentially, the HEMS needs to be able to form and solve a linear program to plan optimal consumption. The key result of our paper is that, under this model and assuming that consumers are using HEMS to control their devices, the grid can deploy pre-announced, time-varying energy prices to accomplish near optimal demand response under a mean-field limit.
The use of time-varying energy prices, also known as time-of-use (TOU) pricing, is one of the oldest ideas for demand response, and discussions of TOU pricing go back to the 19th century (Hausman and Neufeld, 1984). Several recent studies of TOU pricing are skeptical about their potential for meaningful demand response (Schittekatte et al., 2024). Many of these limitations, however, are tied to specifics of existing TOU designs. Most TOU programs that have been deployed to date have involved simple heuristics (e.g., dividing each day into “daytime”, “evening” and “night” and using constant electricity prices within each of these 3 periods), and then relied on consumers to figure out themselves how to respond to these price signals (Cosmo et al., 2014). There are limits to how much such simple TOU programs can shift demand. First of all, they place considerable attention burden on consumers, and even when consumers do attempt to respond to these TOU prices they will generally not have the sophistication to do so optimally (Allcott, 2011a). Second, if consumers were to procure technological devices that enabled them to actually best-respond to such simple TOU prices, new unstable and/or undesirable demand patterns may emerge (Sunar et al., 2024): For example, if everyone times their dishwasher to go on right after high “evening” prices turn to lower “night” prices, then this will create a new demand spike the grid will need to address.
Meanwhile, in the academic literature, there has been considerable interest in deriving optimal time-varying prices under generic choice models where each consumer has a (smooth, concave) welfare function capturing their benefits from energy use (Datchanamoorthy et al., 2011; Yang et al., 2012; Peura and Bunn, 2015; Zhou et al., 2017; Adelman and Uçkun, 2019). Results in this space are of considerable conceptual interest (see the related work section for further discussion); however, there are obstacles to putting the proposed algorithms directly to use in practice. Asking consumers to explicitly reveal their welfare functions would require them to solve a complicated reasoning task, and is thus likely not realistic in practice (Friedman and Oren, 1995; Hobman et al., 2016). Additionally, reduced-form calibration of optimal pricing models requires measuring consumer demand elasticities that may be hard to measure (Mnatsakanyan and Kennedy, 2014).
The modeling assumptions we take in this paper, i.e., with indifference sets to capture flexibility and HEMS to automatically control devices, considerably help alleviate issues with existing time-of-use pricing approaches. We believe it to be far more practical and realistic to ask a consumer to set hard constraints on desired outcomes than to elicit a complete welfare function; for example, it’s easier for a consumer to say they’ll accept indoor temperature between 21–24°C during the day and 20–22°C at night than to specify exactly how their experienced welfare depends on temperature over time. With HEMS, consumers can program their preferences once and avoid actively responding to price signals, all while preserving privacy by not sharing sensitive information with third parties. [^10] As discussed further below, our learning results allow HEMS to operate under a one-way communication model where the HEMS receives prices from the grid and reacts to them, but does not send any information in response (other than the actual energy consumption). Finally, automation allows the grid to use prices that change frequently / gradually enough to avoid derived instability around price changes.
These assumptions, however, introduce novel analytical challenges. Under our indifference set model, consumer demand typically varies discontinuously with prices, rendering many technical tools from existing literature inapplicable (e.g., Adelman and Uçkun, 2019; Yang et al., 2012; Zhou et al., 2017). The bulk of this paper demonstrates how classical results on convex analysis over random sets—building on the foundational work of Aumann (1965) and Artstein and Vitale (1975)—can be leveraged to address these challenges and establish the validity of our approach in large markets. Specifically, we show the following:
@@
    We start by comparing our proposed pricing mechanism to a “direct control mechanism” where the HEMS delegates full control of all user devices to the grid, thus enabling solution of a single global optimization problem that guarantees first-best demand response w.r.t. indifference-sets. We find that, under our model for demand response, the pricing mechanism described above is generally not optimal relative to the direct control mechanism, i.e., that a grid operator with direct control of all consumer devices can achieve better demand response (while respecting user indifference sets) that an operator who can only control demand via dynamic prices. [^11] This finding may seem surprising at first glance, as there are numerous welfare theorems in economics demonstrating that price-based mechanisms can achieve optimal welfare in various systems (Mas-Colell et al., 1995). Our indifference-set model, however, is inconsistent with a key assumption of these welfare theorems, often referred to as local non-satiation. Local non-satiation implies that consumer demand moves smoothly with prices—but this is not the case here, as in our setting we only consider hard constraints on the user side.
@@
    Mean-field optimal prices can be computed via a first-order algorithm that only requires querying HEMS about planned consumption profiles at different price levels, as these consumption profiles give information about a dual cost function. This allows for the straightforward application of first-order optimization methods to find the optimal price using only market-level data, without the grid needing to directly query for individual user preferences.
Finally, we demonstrate practical promise of our pricing approach on a grid-simulation motivated by pre-cooling with data from Phoenix, Arizona; we use OpenDSS to solve the relevant power-flow equations. As seen in Figure 1, renewable production is heavily concentrated during daytime hours. With flat rates, this results in the notorious “duck curve” with low (even negative) net demand during the day that then quickly ramps up in the evening. [^12] The ramp-up period forms the neck of the “duck” and the evening peak demand its head (Denholm et al., 2015). In contrast, our proposed algorithm shifts demand earlier into the day when solar resources are plentiful, thus enabling it to reduce load in the evening hours. We give further details on this experiment in Section 5.
@@ 1.1 Related work
There is a rich literature on using time-varying prices for demand response (Adelman and Uçkun, 2019; Datchanamoorthy et al., 2011; Peura and Bunn, 2015; Schittekatte et al., 2024; Uckun, 2012; Yang et al., 2012; Zhou et al., 2017, 2019). Formal studies of such pricing-based mechanisms are typically framed as Stackelberg games with a two-level leader-follower structure (Adelman and Uçkun, 2019; Yang et al., 2012; Zhou et al., 2017, 2019): First the grid sets prices, and then consumers respond to these prices. Our indifference-set approach is a special case of such a Stackelberg games, where lower-level optimization involves consumers choosing their energy consumption to minimize costs while remaining within their indifference sets (i.e., the consumer preferences are instantiated as hard constraints on allowed consumption patterns). We note that Stackleberg games are also often used for the more abstract task of policy learning with strategic agents (Chen et al., 2020; Munro, 2025; Perdomo et al., 2020; Sahoo and Wager, 2022); and our paper can be seen as a part of this broader literature that exploits specific structure of consumer energy price response.
The major difference between our paper and existing work is that we model consumer preferences using constraints (as captured by indifference sets) rather than using smooth welfare functions—and as a result of that this requires us to develop entirely different technical notions based on convex analysis on random sets to study relevant system properties (as existing analytic strategies fundamentally rely on the smoothness of welfare functions). As discussed above, we believe our indifference-set approach to be substantially more tractable from a practical implementation point of view.
@@
One recent and representative paper from the existing literature is Adelman and Uçkun (2019), who study dynamic pricing in a setting that includes current energy demand, appliance status, operation times, global variables such as weather conditions, among many others. They assume that each consumer’s welfare depends on their energy usage in a differentiable way, and that consumers choose their energy consumption to optimize welfare in response to set prices (e.g., using a HEMS). Given this setup, the authors define the social planner’s welfare function as the sum of all consumers’ welfare minus the grid’s generation cost, and seek prices that optimize the social planner’s welfare under the posited consumer behavior model (the consumers choose how much to consume in response to prices, and the grid must then meet this demand). Their main result is that, as the system size grows large, the optimal price mechanism in their model is one that perfectly externalizes the grid’s incremental production costs to the consumers, i.e., in equilibrium, energy is priced to match the marginal cost of production. The authors propose finding prices that maximize the social planner’s welfare via a tempered fixed-point iteration algorithm.
@@
Like Adelman and Uçkun (2019), we consider a model where—at least in a qualitative sense—prices are set in the view of mutual benefit of consumers and producers. However, unlike Adelman and Uçkun (2019), we do not rely on properties of the consumer welfare function, and do not posit a social planner who is able to trade-off consumer and producer welfare on an additive scale. Then, several mathematical tools used by Adelman and Uçkun (2019)—such as the implicit function theorem to capture consumer response to prices—are no longer applicable in our model where consumer preferences are modeled using indifference set, and so we cannot use analytic strategies from their paper to study our setting.
@@
In another recent line of work, Zhou et al. (2017) consider demand-response programs that employ price incentives, where each consumer aims to minimize their own cost function. This cost function is composed of two parts: One that reflects power consumption and remains independent of price incentives (e.g., minimizing deviation from a nominal consumption level), and another that accounts for the expenses incurred due to price incentives. Building on this framework, Zhou et al. (2019) extends the analysis to incorporate discrete decision variables and dynamically coupled devices over time, such as HVAC systems. Their approach shares similarities with ours, particularly in modeling consumer objectives as minimizing cost functions and solving a social-welfare optimization problem (the direct method in our setup) to determine optimal pricing strategies. However, the authors fundamentally rely on an assumption that consumers set their power consumption by minimizing strongly convex cost functions whereas the cost functions implied by our indifference-set approach are not strongly convex—thus again making their formal results non-applicable to our setting. We also note that the lack of strong convexity in our model results in different qualitative system behavior: In the setting of Zhou et al. (2017) the authors find pricing to always achieve optimal demand response, whereas in our model pricing is not optimal in finite systems (and the optimality gap only closes in a large-sample limit).
@@
Finally, there have also been a number of recent papers studying incentive- rather than price-based mechanisms for demand response. Agrawal and Yücel (2022), Daniels and Lobel (2014), and Webb et al. (2019) explore direct load control contracts, where utility companies manage customer loads, typically to reduce consumption as a form of demand response (in contrast, in the pricing algorithm we evaluate in our experiments, the grid has no explicit control over users’ consumption; instead, HEMS selects optimal consumption profiles that meet users’ needs while minimizing costs). Fattahi et al. (2023) formulates a stochastic dynamic program to design optimal contracts and develop approximation methods to achieve near-optimal solutions for energy consumption (they avoid explicitly referring to these contracts as demand response due to their complexity and the extended nature of the control periods). Allcott (2011b) studies the effect of “home energy reports” that compare a consumer’s energy use to that of their neighbors, and finds that such programs can meaningfully shift behavior.
Our analytic results rely on random set theory to characterize consumer flexibility in demand response programs. Random sets have a rich history in applied probability, going back to Aumann (1965). In particular, the strong laws of large numbers for random of Artstein and Vitale (1975) plays a key role in our analysis. One key early application of these methods was to show existence of competitive equilibria in markets with a continuum of agents who may have non-concave utility functions (Aumann, 1966; Hildenbrand, 1974). More recently, random-set methods have been used in econometrics to study inference in partially identified models (Beresteanu and Molinari, 2008; Beresteanu et al., 2012; Kaido and Santos, 2014). Textbook treatments of random set theory are given in Molchanov (2005) and Artstein-Avidan et al. (2015).
Our results on learning-to-price only assume that the grid can observe energy consumption levels under set prices; and in particular do not require HEMS to share any (potentially sensitive) private information (e.g., information on consumer preferences that underlie its consumption choices). As such, our learning problem can be seen as a bandit problem with a continuous, high-dimensional action space (Ferreira et al., 2018). In general, bandit problems with continuous action spaces do not allow for last learning (Shamir, 2013). However, in our setting, we show that the grid can use mean-field properties of the system to effectively infer relevant gradients from raw consumption data, thus enabling faster convergence properties (Wager and Xu, 2021).
The remainder of the paper is structured as follows. Section 2 introduces the optimal demand response problem and formalizes the indifference set model. Section 3 establishes the mean-field optimality of pricing mechanisms and characterizes their asymptotic performance. Section 4 presents a first-order algorithm to compute optimal prices using only market-level data. Section 5 demonstrates the effectiveness of our approach through a case study on pre-cooling in Phoenix, Arizona. Finally, Section 6 discusses broader implications and concludes.